# A Generalized Analytical Framework for the Nonlinear Best-Worst Method

### v1

@@ 2.3.2 A Secondary Objective Function to Obtain a Unique Weight Set
if $a_{bi}\times a_{iw}\leq\tilde{a}_{bw}^{*}$ and the optimal objective value of minimization problem
@@
if $a_{bi}\times a_{iw}>\tilde{a}_{bw}^{*}$. Furthermore, [47, Theorem 6] states that the PCS
[…]
where $\tilde{a}_{bw}^{*}$ is as defined in equation (18) and
@@
is the only optimally modified PCS. Now, from this PCS, the best optimal weight set can be obtained using Theorem 1.
@@ 3 Research Gap
Although the analytical framework proposed by Wu et al. [47] significantly improves the model’s efficiency, some research gaps remain unaddressed.
(i) Incompatibility of the framework with some scales:
The framework is fully compatible with the Saaty scale but proves incompatible with some alternative scales, including the Salo-Hämäläinen scale, the Lootsma scale and the Donegan-Dodd-McMaster scale.
@@
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equation (13), we get $\epsilon_{2}=0.4056$, $\epsilon_{3}=0.3944$ and $\epsilon_{4}=0.5363$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=0.5163$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})<a_{15}-\epsilon_{4}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4}=0.5363$. The optimal interval-weights computed using equation (20) are given in Table 3.
Now, using equation (18a), we have $\tilde{a}_{bw}^{*}=8.4638$. By equation (24), we get $\eta_{2}^{*}=0.5038$, $\eta_{3}^{*}=0.5481$ and $\eta_{4}^{*}=0.5363$. From equation (23), it follows that the best optimally modified PCS is
@@
The best optimal weight set calculated using Theorem 1 is presented in Table 3.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the derived best optimal weight set should not be regarded as an optimal weight set.
We calculated the actual optimal interval-weights and the best optimally modified PCS using MATLAB. These optimal interval-weights are shown in Table 3. The best optimally modified PCS is
@@
Using Theorem 1, we derived the best optimal weight set, which is also provided in Table 3.
@@
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equation (13), we get $\epsilon_{2}=1.6054$, $\epsilon_{3}=1.4764$ and $\epsilon_{4}=1.7228$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=1.7882$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{2})\times(a_{45}+\epsilon_{2})<a_{15}+\epsilon_{2}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,2}=1.7882$. The optimal interval-weights computed using equation (20) are given in Table 4.
Now, using equation (18a), we have $\tilde{a}_{bw}^{*}=14.7841$. By equation (24), we get $\eta_{2}^{*}=1.7882$, $\eta_{3}^{*}=1.8118$ and $\eta_{4}^{*}=1.7882$. From equation (23), it follows that the best optimally modified PCS is
@@
The best optimal weight set calculated using Theorem 1 is presented in Table 4.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the derived best optimal weight set should not be regarded as an optimal weight set.
We calculated the actual optimal interval-weights and the best optimally modified PCS using MATLAB. These optimal interval-weights are shown in Table 4. The best optimally modified PCS is
@@
Using Theorem 1, we derived the best optimal weight set, which is also provided in Table 4.
@@
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equation (13), we get $\epsilon_{2}=0.6959$, $\epsilon_{3}=0.6781$ and $\epsilon_{4}=0.8086$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=0.8825$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{2})\times(a_{45}+\epsilon_{2})<a_{15}+\epsilon_{2}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,2}=0.8825$. The optimal interval-weights computed using equation (20) are given in Table 5.
Now, using equation (18a), we have $\tilde{a}_{bw}^{*}=5.3640$. By equation (24), we get $\eta_{2}^{*}=0.8825$, $\eta_{3}^{*}=0.9129$ and $\eta_{4}^{*}=0.8825$. From equation (23), it follows that the best optimally modified PCS is
@@
The best optimal weight set calculated using Theorem 1 is presented in Table 5.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the derived best optimal weight set should not be regarded as an optimal weight set.
We calculated the actual optimal interval-weights and the best optimally modified PCS using MATLAB. These optimal interval-weights are shown in Table 5. The best optimally modified PCS is
@@
Using Theorem 1, we derived the best optimal weight set, which is also provided in Table 5.
@@
(ii) Incompatibility of the framework with the Saaty scale in the presence of multiple DMs:
@@
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$ and $D_{3}=\phi$. By equation (13), we get $\epsilon_{2}=0.4355$, $\epsilon_{3}=0.4401$ and $\epsilon_{4}=0.4865$. This gives $i_{1}=4$ and $i_{2}=3$. Now, by equation (16), we get $\epsilon_{4,3}=0.5458$. Note that $(a_{13}-\epsilon_{4})\times(a_{35}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{3})\times(a_{45}+\epsilon_{3})<a_{15}+\epsilon_{3}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,3}=0.5458$. The optimal interval-weights computed using equation (20) are given in Table 6.
Now, using equation (18a), we have $\tilde{a}_{bw}^{*}=6.8230$. By equation (24), we get $\eta_{2}^{*}=0.5502$, $\eta_{3}^{*}=0.5458$ and $\eta_{4}^{*}=0.5458$. From equation (23), it follows that the best optimally modified PCS is
@@
The best optimal weight set calculated using Theorem 1 is presented in Table 6.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{2}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the derived best optimal weight set should not be regarded as an optimal weight set.
We calculated the actual optimal interval-weights and the best optimally modified PCS using MATLAB. These optimal interval-weights are shown in Table 6. The best optimally modified PCS is
@@
Using Theorem 1, we derived the best optimal weight set, which is also provided in Table 6.
@@
The CI for the Saaty scale, presented in Table 2, does not serve as an upper bound for the set of $\epsilon^{*}$ corresponding to the PCSs having the given $a_{bw}$, and thus, is not well-defined.
@@
In instances where multiple criteria are equally qualified as the best (or worst), the conventional approach involves arbitrarily selecting one as the best (or worst). However, the chosen one might get a different weight than the others even though they are of equal preference.
@@
The interval-weights, the best weight set and $\epsilon^{*}$ are given in Table 8.
Note that for $c_{1}$ and $c_{2}$, neither the interval-weights nor the optimal weights coincide, despite both criteria being of equal preferences.
@@
4 A Generalized Analytical Framework for the Nonlinear Best-Worst Method
In this section, we propose a generalized analytical framework that is compatible with any scale and any number of DMs. We derive a formula for CI and an analytical expression for CR. Furthermore, we modify the original optimization model to ensure equal weights for criteria with identical preferences when multiple best/worst criteria are present.
@@ 4.1 Calculation of Weights
4.1.1 Calculation of Weights in the Presence of Unique Best and Worst Criterion
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\eta^{*}$ be the optimal objective value of problem (7). Then $\epsilon_{i}\leq\eta^{*}$ and $\epsilon_{i,j}\leq\eta^{*}$ for all $i,j\in D$.
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, and let $\eta^{*}$ be the optimal objective value of problem (7). Then the following statements hold.
@@
    Here, we get $(a_{bi}+a_{iw})\leq\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}$. Therefore,
    $\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\geq 0,$
    and thus,
@@
    This gives
    $|\tilde{a}_{bi}-a_{bi}|=\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\leq\epsilon_{i}\leq\epsilon_{i_{0}}.$
    Similarly, we get $|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0}}$.
@@
    Here, we get $\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}<a_{bi}+a_{iw}$. Therefore,
    $\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\geq 0,$
    and thus,
@@
    This gives
    $|\tilde{a}_{bi}-a_{bi}|=\frac{a_{bi}+a_{iw}-\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}}{2}\leq\epsilon_{i,i_{0}}\leq\epsilon_{i_{0}}.$
    Similarly, we get $|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0}}$.
@@
    Here, we get $(a_{bi}+a_{iw})\leq\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}$. So,
    $\frac{-(a_{bi}+a_{iw})+\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}}{2}\geq 0,$
    and thus,
@@
From equation 2, it follows that
$\begin{split}\epsilon^{*}=\eta^{*}&=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}\\ &=\max\left\{\bigg{|}\frac{a_{bi}+a_{iw}+1-\sqrt{(a_{bi}+a_{iw}+1)^{2}-4(a_{bi}\times a_{iw}-a_{bw})}}{2}\bigg{|},\right.\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\bigg{|}\frac{a_{bi}\times a_{iw}-a_{bj}\times a_{jw}}{a_{bi}+a_{iw}+a_{bj}+a_{jw}}\bigg{|}:i,j\in D\right\},\end{split}$ (33)
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equation (13) and (16) respectively, let $\eta^{*}$ be the optimal objective value of problem (7), and let $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ be an optimally modified PCS. Then
@@
First, assume that $\eta^{*}=\epsilon_{i_{0}}\text{ for some }i_{0}\in D_{1}$. As discussed in Proposition 1, we get $0\leq\eta_{bi_{0}},\eta_{i_{0}w},\eta_{bw}\leq\eta^{*}=\epsilon_{i_{0}}$ such that $\tilde{a}_{bi_{0}}^{*}\in\{a_{bi_{0}}-\eta_{bi_{0}},a_{bi_{0}}+\eta_{bi_{0}}\}$, $\tilde{a}_{i_{0}w}^{*}\in\{a_{i_{0}w}-\eta_{i_{0}w},a_{i_{0}w}+\eta_{i_{0}w}\}$ and $\tilde{a}_{bw}^{*}\in\{a_{bw}-\eta_{bw},a_{bw}+\eta_{bw}\}$. Since $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ is consistent, we have $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bw}^{*}$. Now, it is easy to observe that if any one of $\eta_{bi_{0}}$, $\eta_{i_{0}w}$ or $\eta_{bw}$ is strictly less than $\epsilon_{i_{0}}$, then at least one of the remaining two must be strictly greater than $\epsilon_{i_{0}}$, which is not possible. This gives $\eta_{bi_{0}}=\eta_{i_{0}w}=\eta_{bw}=\epsilon_{i_{0}}$. It is clear that the only combination among the eight possible values of $(\tilde{a}_{bi_{0}}^{*},\tilde{a}_{i_{0}w}^{*},\tilde{a}_{bw}^{*})$ that satisfies $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bw}^{*}$ is $(a_{bi_{0}}+\epsilon_{i_{0}},a_{i_{0}w}+\epsilon_{i_{0}},a_{bw}-\epsilon_{i_{0}})$. Therefore, $\tilde{a}_{bw}^{*}=a_{bw}-\epsilon_{i_{0}}$. The result follows by an analogous argument if $\eta^{*}=\epsilon_{j_{0}}\text{ for some }j_{0}\in D_{2}$.

### v2

@@ 2.3.2 A Secondary Objective Function
if $a_{bi}\times a_{iw}\leq\tilde{a}_{bw}^{*}$ and the optimal objective value of problem
@@
if $a_{bi}\times a_{iw}>\tilde{a}_{bw}^{*}$, where $\tilde{a}_{bw}^{*}$ is as in equation (18a). Furthermore, [59, Theorem 6] states that the PCS
$\begin{split}&\tilde{a}_{bi}^{*}=\begin{cases}a_{bi}+\eta_{i}^{*}\quad\text{if }a_{bi}\times a_{iw}\leq\tilde{a}_{bw}^{*},\\ a_{bi}-\eta_{i}^{*}\quad\text{if }a_{bi}\times a_{iw}>\tilde{a}_{bw}^{*},\end{cases}\\ &\tilde{a}_{iw}^{*}=\begin{cases}a_{iw}+\eta_{i}^{*}\quad\text{if }a_{bi}\times a_{iw}\leq\tilde{a}_{bw}^{*},\\ a_{iw}-\eta_{i}^{*}\quad\text{if }a_{bi}\times a_{iw}>\tilde{a}_{bw}^{*},\text{ where}\end{cases}\end{split}$ (23)
@@
is the only best optimally modified PCS. Using this PCS, the best optimal weight set is obtained using equation (8).
@@ 3 Research Gap
Despite the wide applicability of the nonlinear BWM, several research gaps remain to be addressed.
(i) Incompatibility of the existing analytical framework with some scales:
The existing analytical approach is fully compatible with the Saaty scale but proves incompatible with some other scales such as the Salo-Hämäläinen scale, the Lootsma scale, and the Donegan-Dodd-McMaster scale.
@@
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$, and $D_{3}=\emptyset$. By equation (13), we get $\epsilon_{2}=0.4056$, $\epsilon_{3}=0.3944$, and $\epsilon_{4}=0.5363$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=0.5163$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})<a_{15}-\epsilon_{4}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4}=0.5363$. The optimal interval-weights computed using equation (20) are given in Table 3.
Now, from equation (18a), we have $\tilde{a}_{bw}^{*}=8.4638$. By equation (24), we get $\eta_{2}^{*}=0.5038$, $\eta_{3}^{*}=0.5481$, and $\eta_{4}^{*}=0.5363$. From equation (23), it follows that the best optimally modified PCS is
@@
The best optimal weight set calculated using equation (8) is presented in Table 3.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered as an optimally modified PCS. Consequently, the obtained best optimal weight set should not even be regarded as an optimal weight set.
We calculated the actual $\epsilon^{*}$, optimal interval-weights and the best optimally modified PCS using MATLAB. These $\epsilon^{*}$ and optimal interval-weights are given in Table 3. The best optimally modified PCS is
@@
Using equation (8), we derived the best optimal weight set, which is also provided in Table 3.
@@
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$, and $D_{3}=\emptyset$. By equation (13), we get $\epsilon_{2}=1.6054$, $\epsilon_{3}=1.4764$, and $\epsilon_{4}=1.7228$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=1.7882$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{2})\times(a_{45}+\epsilon_{2})<a_{15}+\epsilon_{2}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,2}=1.7882$. The optimal interval-weights computed using equation (20) are given in Table 4.
Now, from equation (18a), we have $\tilde{a}_{bw}^{*}=14.7841$. By equation (24), we get $\eta_{2}^{*}=1.7882$, $\eta_{3}^{*}=1.8118$, and $\eta_{4}^{*}=1.7882$. From equation (23), it follows that the best optimally modified PCS is
@@
The best optimal weight set calculated using equation (8) is presented in Table 4.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the obtained best optimal weight set should not even be regarded as an optimal weight set.
We calculated the actual $\epsilon^{*}$, optimal interval-weights and the best optimally modified PCS using MATLAB. These $\epsilon^{*}$ and optimal interval-weights are given in Table 4. The best optimally modified PCS is
@@
Using equation (8), we derived the best optimal weight set, which is also provided in Table 4.
@@
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$, and $D_{3}=\emptyset$. By equation (13), we get $\epsilon_{2}=0.6959$, $\epsilon_{3}=0.6781$, and $\epsilon_{4}=0.8086$. This gives $i_{1}=4$ and $i_{2}=2$. Now, by equation (16), we get $\epsilon_{4,2}=0.8825$. Note that $(a_{12}-\epsilon_{4})\times(a_{25}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{2})\times(a_{45}+\epsilon_{2})<a_{15}+\epsilon_{2}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,2}=0.8825$. The optimal interval-weights computed using equation (20) are given in Table 5.
Now, from equation (18a), we have $\tilde{a}_{bw}^{*}=5.3640$. By equation (24), we get $\eta_{2}^{*}=0.8825$, $\eta_{3}^{*}=0.9129$, and $\eta_{4}^{*}=0.8825$. From equation (23), it follows that the best optimally modified PCS is
@@
The best optimal weight set calculated using equation (8) is presented in Table 5.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{3}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the obtained best optimal weight set should not even be regarded as an optimal weight set.
We calculated the actual $\epsilon^{*}$, optimal interval-weights and the best optimally modified PCS using MATLAB. These $\epsilon^{*}$ and optimal interval-weights are given in Table 5. The best optimally modified PCS is
@@
Using equation (8), we derived the best optimal weight set, which is also provided in Table 5.
@@
(ii) Incompatibility of the existing framework in the presence of multiple DMs:
@@
Here, $D_{1}=\{4\}$, $D_{2}=\{2,3\}$, and $D_{3}=\emptyset$. By equation (13), we get $\epsilon_{2}=0.4355$, $\epsilon_{3}=0.4401$, and $\epsilon_{4}=0.4865$. This gives $i_{1}=4$ and $i_{2}=3$. Now, by equation (16), we get $\epsilon_{4,3}=0.5458$. Note that $(a_{13}-\epsilon_{4})\times(a_{35}-\epsilon_{4})>a_{15}-\epsilon_{4}$ and $(a_{14}+\epsilon_{3})\times(a_{45}+\epsilon_{3})<a_{15}+\epsilon_{3}$. So, equation (17) gives $\epsilon^{*}=\epsilon_{4,3}=0.5458$. The optimal interval-weights computed using equation (20) are given in Table 6.
Now, from equation (18a), we have $\tilde{a}_{bw}^{*}=6.8230$. By equation (24), we get $\eta_{2}^{*}=0.5502$, $\eta_{3}^{*}=0.5458$, and $\eta_{4}^{*}=0.5458$. From equation (23), it follows that the best optimally modified PCS is
@@
The best optimal weight set calculated using equation (8) is presented in Table 6.
It is important to note that the calculated interval-weights are not well-defined, as the lower bound of each interval-weight exceeds the upper bound. Additionally, the value of $\eta_{2}^{*}$ is higher than $\epsilon^{*}$, meaning the resulting best optimally modified PCS should not even be considered an optimally modified PCS. Consequently, the obtained best optimal weight set should not even be regarded as an optimal weight set.
We calculated the actual $\epsilon^{*}$, optimal interval-weights and the best optimally modified PCS using MATLAB. These $\epsilon^{*}$ and optimal interval-weights are given in Table 6. The best optimally modified PCS is
@@
Using equation (8), we derived the best optimal weight set, which is also provided in Table 6.
@@
The values of CI for the Saaty scale, presented in Table 2, do not serve as an upper bound for the set of $\epsilon^{*}$ corresponding to the PCSs having the given $a_{bw}$, and thus, is not well-defined.
@@
In instances where multiple criteria are equally qualified as the best (or worst), the conventional approach involves arbitrarily selecting one as the best (or worst) for computational purposes. However, the chosen one might get a different weight than the others even though they are of equal preference.
@@
The interval-weights, the best optimal weight set and $\epsilon^{*}$ are given in Table 8.
Note that for $c_{1}$ and $c_{2}$, neither the interval-weights nor the best weights coincide, despite both criteria being of equal preferences.
@@
4 A Generalized Analytical Framework for the Nonlinear BWM
In this section, we propose an analytical approach compatible with any scale and any number of DMs. We also modify the original optimization model to ensure equal weights for all best and worst criteria in cases where multiple best or worst criteria are present. Furthermore, we derive formulas for CI and CR.
@@ 4.1 Calculation of Weights
4.1.1 Unique Best and Worst Criterion
It is easy to verify that equations (18a) and (20) hold in general, i.e., they are valid for any scale and any number of DMs. Therefore, to determine the optimal interval-weights, it suffices to derive the general expressions for $\epsilon^{*}$ and $\tilde{a}_{bw}^{*}$. In the following, our objective is to compute these values. We also derive the best optimally modified PCS that leads to the best optimal weight set.
@@ 4.1.1 Unique Best and Worst Criterion
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equations (13) and (16) respectively, and let $\eta^{*}$ be the optimal objective value of problem (7). Then $\epsilon_{i}\leq\eta^{*}$ and $\epsilon_{i,j}\leq\eta^{*}$ for all $i,j\in D$.
@@
Proposition 1 establishes that $\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}$ serves as a lower bound for $\eta^{*}$ (and consequently, for $\epsilon^{*}$).
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equations (13) and (16) respectively, and let $\eta^{*}$ be the optimal objective value of problem (7). Then the following statements hold.
@@
    Here, we get $(a_{bi}+a_{iw})\leq\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}$. Therefore, -(abi+aiw)+(abi+aiw)2-4×abi×aiw+4(abw-ϵi0)2≥0, and thus,
@@
    This gives —~a_bi-a_bi—=-(abi+aiw)+(abi+aiw)2-4×abi×aiw+4(abw-ϵi0)2≤ϵ_i≤ϵ_i_0. Similarly, we get $|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0}}$.
@@
    Here, we get $\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bw}-\epsilon_{i_{0}})}<a_{bi}+a_{iw}$. Therefore, abi+aiw-(abi+aiw)2-4×abi×aiw+4(abw-ϵi0)2≥0, and thus,
@@
    This gives —~a_bi-a_bi—=abi+aiw-(abi+aiw)2-4×abi×aiw+4(abw-ϵi0)2≤ϵ_i,i_0≤ϵ_i_0. Similarly, we get $|\tilde{a}_{bi}-a_{bi}|\leq\epsilon_{i_{0}}$.
@@
    Here, we get $(a_{bi}+a_{iw})\leq\sqrt{(a_{bi}+a_{iw})^{2}-4\times a_{bi}\times a_{iw}+4(a_{bi_{0}}+\epsilon_{i_{0},j_{0}})\times(a_{i_{0}w}+\epsilon_{i_{0},j_{0}})}$. So, -(abi+aiw)+(abi+aiw)2-4×abi×aiw+4(abi0+ϵi0,j0)×(ai0w+ϵi0,j0)2≥0, and thus,
@@
Theorem 2, along with the facts that $\epsilon_{i}=0$ for $i\in D_{3}$, $\epsilon_{i,j}=\epsilon_{j,i}$ for all $i,j\in D$, and that $\epsilon_{i,j}<\max\{\epsilon_{i},\epsilon_{j}\}$ whenever $i,j\in D_{1}\cup D_{3}$ or $i,j\in D_{2}\cup D_{3}$, gives
$\begin{split}\epsilon^{*}=\eta^{*}&=\max\{\epsilon_{i},\epsilon_{i,j}:i,j\in D\}=\max\{\epsilon_{i},\epsilon_{j,k}:i\in D_{1}\cup D_{2},j\in D_{1},k\in D_{2}\}\\ &=\max\left\{\bigg|\frac{a_{bi}+a_{iw}+1-\sqrt{(a_{bi}+a_{iw}+1)^{2}-4(a_{bi}\times a_{iw}-a_{bw})}}{2}\bigg|,\right.\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\ \ \left.\bigg|\frac{a_{bi}\times a_{iw}-a_{bj}\times a_{jw}}{a_{bi}+a_{iw}+a_{bj}+a_{jw}}\bigg|:i,j\in D\right\},\\ &=\max\left\{\bigg|\frac{a_{bi}+a_{iw}+1-\sqrt{(a_{bi}+a_{iw}+1)^{2}-4(a_{bi}\times a_{iw}-a_{bw})}}{2}\bigg|,\right.\\ &\quad\quad\quad\quad\quad\quad\quad\quad\ \ \ \left.\frac{a_{bk}\times a_{kw}-a_{bj}\times a_{jw}}{a_{bj}+a_{jw}+a_{bk}+a_{kw}}:i\in D_{1}\cup D_{2},j\in D_{1},k\in D_{2}\right\},\end{split}$ (33)
@@
Let $\epsilon_{i}$ and $\epsilon_{i,j}$ be as in equations (13) and (16) respectively, let $\eta^{*}$ be the optimal objective value of problem (7), and let $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ be an optimally modified PCS. Then
@@
First, assume that $\eta^{*}=\epsilon_{i_{0}}\text{ for some }i_{0}\in D_{1}$. As discussed in Proposition 1, we get $0\leq\eta_{bi_{0}},\eta_{i_{0}w},\eta_{bw}\leq\eta^{*}=\epsilon_{i_{0}}$ such that $\tilde{a}_{bi_{0}}^{*}\in\{a_{bi_{0}}-\eta_{bi_{0}},a_{bi_{0}}+\eta_{bi_{0}}\}$, $\tilde{a}_{i_{0}w}^{*}\in\{a_{i_{0}w}-\eta_{i_{0}w},a_{i_{0}w}+\eta_{i_{0}w}\}$, and $\tilde{a}_{bw}^{*}\in\{a_{bw}-\eta_{bw},a_{bw}+\eta_{bw}\}$. Since $(\tilde{A}_{b}^{*},\tilde{A}_{w}^{*})$ is consistent, we have $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bw}^{*}$. Now, it is easy to observe that if any one of $\eta_{bi_{0}}$, $\eta_{i_{0}w}$ or $\eta_{bw}$ is strictly less than $\epsilon_{i_{0}}$, then at least one of the remaining two must be strictly greater than $\epsilon_{i_{0}}$, which is not possible. This gives $\eta_{bi_{0}}=\eta_{i_{0}w}=\eta_{bw}=\epsilon_{i_{0}}$. It is clear that the only combination among the eight possible values of $(\tilde{a}_{bi_{0}}^{*},\tilde{a}_{i_{0}w}^{*},\tilde{a}_{bw}^{*})$ that satisfies $\tilde{a}_{bi_{0}}^{*}\times\tilde{a}_{i_{0}w}^{*}=\tilde{a}_{bw}^{*}$ is $(a_{bi_{0}}+\epsilon_{i_{0}},a_{i_{0}w}+\epsilon_{i_{0}},a_{bw}-\epsilon_{i_{0}})$. Therefore, $\tilde{a}_{bw}^{*}=a_{bw}-\epsilon_{i_{0}}$. The result follows by an analogous argument if $\eta^{*}=\epsilon_{j_{0}}\text{ for some }j_{0}\in D_{2}$.